OpenStudy (anonymous):

I need to change ƒ(x)=(x-3)^2 to standard form. I don't even know where to start thanks in advance!

OpenStudy (acxbox22):

f(x)=(x-3)^2 f(x)=(x-3)(x-3) f(x)=x^2-3x-3x+9 f(x)=x^2-6x+9

OpenStudy (loser66):

thank you!!:)

OpenStudy (acxbox22):

np

OpenStudy (loser66):

lol